Vin Diesel owns the Sheffield Barber Shop. He employs 7 barbers and pays each a base rate of $1,570 per month. One of the barbers serves as the manager and receives an extra $540 per month. In addition to the base rate, each barber also receives a commission of $4.50 per haircut.

Other costs are as follows.
Advertising $310 per month
Rent $920 per month
Barber supplies $0.30 per haircut
Utilities $180 per month plus $0.20 per haircut
Magazines $20 per month

Vin currently charges $10.00 per haircut.
Determine the variable costs per haircut and the total monthly fixed costs. (Round variable costs to 2 decimal places, e.g. 2.25.)
Total variable cost per haircut $

Total fixed $

Compute the break-even point in units and dollars. (Round answers to 0 decimal places, e.g. 1,225.)
Break-even point in units

haircuts
Break-even point in dollars $

Determine net income, assuming 3,020 haircuts are given in a month.
Net income / (Loss) $
Answer #1

Total Variable cost = 4.50+0.30+0.20

= 5

Total fixed cost = (7*1570)+540+310+920+180+20

= 12,960

Breakeven point in units = Fixed cost/Contribution margin per unit

= 12,960/(10-5)

= 2592

​​​​​​Breakeven point in dollars = 2592*10

= 25,920

Net income = (3020*5) - 12,960

= 2140

Comment if you face any issues